Hi there, I really need some help on this problem. I have some of it worked out, but I'm confused on the second part:
Let f (x) = 2x2 + 8x – 5 and g(x) = 4x – 1
(a) Find the composite function (f o g)(x) and simplify the results. Show work.
5

(b) Find (f o g)(-1) . Show work.
I'm not sure what I need to do on part b, any help is greatly appreciated!

Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39617)

Evaluate the composition for x=-1.

Answer by stanbon(75887)

Let f (x) = 2x^2 + 8x – 5 and g(x) = 4x – 1
(a) Find the composite function (f o g)(x) and simplify the results. Show work.
fog(x) = f[g(x)] = f[4x-1] = 2(4x-1)^2+8(4x-1)-5
= 2(16x^2 -8x+1)+32x-8-5
----
= 32x^2 - 16x + 2 + 32x - 13
----
= 32x^2 + 16x - 11
-----------------------------
(b) Find (f o g)(-1) = f[g(-1] = f[-4-1] = f[-5]
= 2(-5)^2 + 8(-5) -5
= 50 - 45
---
= 5
-----------
Cheers,
Stan H.
-------------